OpenStudy (anonymous):

Evaluate the following expression for a = 4 and b = -5. a2 - 10 ---------- 2 + b Select one of the options below as your answer: A. -8 B. 6 C. 8 D. -2

OpenStudy (anonymous):

what exactily are you looking for

OpenStudy (anonymous):

answer

OpenStudy (anonymous):

plug in the values so 4^2 = 16-10 = 6 2+-5 = -3 so 6/-3 = -2 -2 is the answer

OpenStudy (anonymous):

D